OpenStudy (bananas1234):

Find the solution to the following equation by transforming it into a perfect square trinomial. x2 – 10x = 39. Is {–3, 13} correct?

OpenStudy (alyssa_xo):

I don't understand that answer. complete the square

OpenStudy (alyssa_xo):

(-5)^2=25 x^2-10x+25

OpenStudy (anonymous):

\[add~5^2~on both~ sides~ then~solve\]

OpenStudy (bananas1234):

These are my options. {–3, 13} {–15, 25} {–20, 30} {59, 69}

OpenStudy (bananas1234):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^2 - 10x = 39 x^2 - 10x + 25 = 39 + 25 ... take half of 10, then square it, and add it to both sides (x-5)^2 = 64 I'll let you finish up